OpenStudy (anonymous):

Find the simple interest.(The rate is an annual rate otherwise noted. Assume 360 days in a year.) Round the nearest cent. p=$1,000, r=5.6%, t=1/3 years

OpenStudy (anonymous):

for simple interest, i = P*r*t) = 1000*.056*1/3) = $18.67

OpenStudy (anonymous):

ty, I foget the formula rule